Consider the function, f x( ) = 2x3 −24x2 −7. Find the intervals of x where f(x) is increasing or decreasing.


1
Expert's answer
2022-03-03T18:01:51-0500

"\frac{df}{dx}=6x^2-48x=6x(x-8)."

"\frac{df}{dx}>0 \to x<0\; or\; x>8".

Thus, "f(x)" increases on "(-\infin,0)\cup(8,\infin)."

"\frac{df}{dx}<0 \to 0<x<8."

Thus, "f(x)" decreases on "(0,8)."